On R^n show that || . ||∞ ≤ || . ||2

.


1
Expert's answer
2020-06-22T18:02:49-0400

Let,"x\in \mathbb{R}^n" such that "x=(x_1,\dots,x_n)" ,thus

"||x||_2=\sqrt{\sum_{k=1}^{n}|x_i|^2}"

And,

"||x||_{\infty}=\max_{1\leq i \leq n}\{|x_i|\}"

Let, WLOG , "||x||_{\infty}=x_k" for some "k\in\{1,\dots,n\}" but also note that


"||x||_{\infty}^2=x_k^2\leq x_1^2+\dots+x_k^2+\dots+x_n^2=||x||_2^2\\
\iff ||x||_{\infty}\leq ||x||_2"

Hence, we are done.